OpenStudy (anonymous):

reduce this radical radical 15 times radical 12

OpenStudy (anonymous):

\[\sqrt{15} * \sqrt{12} = \sqrt{15} * 2\sqrt{3} =2 \sqrt{45} = 2 \sqrt{9*5} = 6\sqrt{5}\]